Light from a laser passes through a pair of slits and forms a pattern on a screen 4
meters from the slits. The slits are 50?m wide and are 0.1mm apart.

a) If the wavelength of the laser is 650 nm, sketch the pattern made on the screen.

b) Calculate the spacing between fringes and the width of the central maximum.

c) A wedge of material is slipped in front of one slit until the central bright fringe disappears (a
dark fringe is produced at the center of the pattern. The thickness of the material in front of the
slit is measured to be 8.45 ?m. What is the index of refraction of the material?
Answer #1

a)

b)

d E

spacing between fringes = \lambda L/d = (650*0.1/50)*10^{-6} = 1.3*10^{-6}

distance between two adjacent dark fringes = width of central maxima = 2*spacing between fringes = 2.6*10^{-6}

c)

 

Optical path difference at

P = S2P - [S1P+